OpenStudy (anonymous):

Alan is twice as old as sue and half as old as Joseph. If the average of all three ages is 14, how old is sue?

terenzreignz (terenzreignz):

This is a strange age-problem :3 But, a word problem nonetheless... the first step in solving word problems is representing values... If we let x be the age of Sue, what is the age of Alan?

OpenStudy (anonymous):

2x

OpenStudy (anonymous):

And Joseph is 1/2x ?

terenzreignz (terenzreignz):

No... Alan's age is 2x, yes, and Alan's age is HALF of Joseph's age... So Joseph must be older... so it's not \(\Large \frac12x\) :P

OpenStudy (anonymous):

S Joseph age is 1/2(2x)

terenzreignz (terenzreignz):

No. Look, let's say Joseph's age is j. And Alan's age, which is 2x is half of j. \[\Large 2x = \frac12j\] So what is j in terms of x?

OpenStudy (anonymous):

4x=j

terenzreignz (terenzreignz):

Better :P So Joseph's age is 4x. Sue = x Alan = 2x Joseph = 4x What's their average?

OpenStudy (anonymous):

14

terenzreignz (terenzreignz):

Yeah, but how do you represent their average?

OpenStudy (anonymous):

X+2x+4x/3=14

terenzreignz (terenzreignz):

Very good :) \[\Large \frac{x+2x+4x}{3}=14\] can you solve for x?

OpenStudy (anonymous):

X=6

terenzreignz (terenzreignz):

And that's Sue's age :D Remember, we let Sue's age be x Good job ^_^

OpenStudy (anonymous):

Ty metal for you

OpenStudy (anonymous):

I hate word problems

terenzreignz (terenzreignz):

Well, they're not gonna love you either way, so you better learn how to DEAL WITH THEM. LOL Signing off now ------------------------------------------ Terence out